OpenStudy (anonymous):

Find the midpoint of the segment whose endpoints are A(2, –3) and B(6, 7).

OpenStudy (anonymous):

wouldnt you do x2-x1/y2-y1?

OpenStudy (anonymous):

(4,2) is the mid point

OpenStudy (anonymous):

[(x1+x2)/2,(y1+y2)/2]

OpenStudy (anonymous):

midpoint of a segment with one end as \[(x_1,y_1)\] and (\[(x_2,y_2)\]=\[\frac{ x_1+x_2 }{ 2 },\frac{ y_1+y_2 }{ 2 }\]

OpenStudy (anonymous):

The midpoint x and y value is the average of the endpoint x and y values. Average the x's, average the y's.

OpenStudy (anonymous):

so according to ur question we have\[\frac{ 2+6 }{ 2 },\frac{ -3+7 }{ 2 }\]=\[4,2\]

OpenStudy (anonymous):

Thank you all!!

OpenStudy (anonymous):

my pleasure!

OpenStudy (anonymous):

thats wat i text

OpenStudy (anonymous):

so u want a medal?

OpenStudy (anonymous):

no i'm good

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

Do math every day.